OpenStudy (anonymous):

9y^2-12y+13=0 solve by factoring or by completing the square

hero (hero):

This quadratic has no real solutions

hero (hero):

It's factorable, it just won't look pretty

OpenStudy (anonymous):

it has to

OpenStudy (anonymous):

i need help don't understand it

hero (hero):

Are all of your signs correct? Double check

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or solve by completing the square

hero (hero):

I'm sorry to hear that. It has no real solutions

OpenStudy (anonymous):

i typed that and i got it wrong. can you solve by completing the square

hero (hero):

You could, but it wouldn't yield a real solution

hero (hero):

Yeah, satellite, help me out here

OpenStudy (anonymous):

complete the square as follows \[9y^2-12y+13=0\] subtract 13 \[9y^2-12y=-13\] divide by 9 \[y^2-\frac{4}{3}y=-\frac{13}{9}\] take half the coefficient of the middle term square and add to both sides to make a perfect square \[(y-\frac{2}{3})^2=-\frac{13}{9}+\frac{4}{9}=-\frac{9}{9}=-1\] take the square root \[y-\frac{2}{3}=\pm \sqrt{-1}=\pm i\] solve for y \[y=\frac{2}{3}\pm i\]

OpenStudy (anonymous):

you can also use the quadratic formula, but you will not be able to factor unless you know the zeros to begin with

hero (hero):

What about myininaya's method?

OpenStudy (anonymous):

thanks that helped alot